Sometimes a person cannot clearly see objects close up or far away. To correct this type of vision, bifocals are often used. The top half of the lens is used to view distant objects and the bottom half of the lens is used to view objects close to the eye. A person can clearly see objects only if they are located between 44 cm and 194 cm away from his eyes. Bifocal lenses are used to correct his vision. What power lens (in diopters) should be used in the top half of the lens to allow him to clearly see distant objects?
Answer #1

Using the lens equation 1/f = 1/p + 1/q

Here as her far point is 194 cm, we need to take objects from infinity and focus them at 194 cm away

Since, The images are virtual and will need a negative by sign convention, because the images are on the same side of the lens as the object therefore

1/f = 1/infinity + 1/-194

f = -194 cm

The power of the lens is the inverse of the focal length, converted to meters.

That is 1/-1.94= -0.515